Tohban Report 2015-02-25

From RHESSI Wiki

Jump to: navigation, search

Tohban Reports
Start Date: 18 Feb 2015
End Date: 25 Feb 2015
Tohban: Milo Buitrago-Casas
Tohban email:
Next Tohban: Milo Buitrago-Casas
List all reports


Solar Activity

Solar activity was very low last week with no M-class flares and some C-class flares. Currently there is only one active region on the disk.

How many GOES flares occurred?

 Flares above B, C, M, X class were     15    10     0     0

And how many of these are listed in the RHESSI flare list?

 Flares above B, C, M, X class were     12     5     0     0

And how many had EXCELLENT coverage?

 Flares above B, C, M, X class were      0     0     0     0

There were RHESSI flares/GOES flares 164 / 25 over the time range 17-Feb-15 24-Feb-15

Memory Management

The SSR was emptying by the end of the daily pass set during the whole week. Decimation was changed from Active/Vigorous to Normal/Vigorous on Feb 19, 2015 at 17:18 UTC. In addition, we changed night time data collection for the front and rear segments to record data during whole eclipses. RHESSI's front and rear segments are recording complete night times since February 19, 2015 at 17:18 UTC.

Spacecraft Status

On February 19, we carried out an experiment to figure it out the data gap problem. Ops deleted out all WLPs passes at the beginning of a pass set that came before the first BGS pass (Since February 19 until February 22). The deleted WLPs passes were compensated by scheduling alternate WLPs views and adding in extra AGO supports.The gap between BGS passes never exceeds 17 hours during the experiment.

It was discovered that WLPs was using a stale ephemeris to track HESSI. The TLE at WLPs was updated on February 19 morning.

Data Gaps

17-Feb-15 00:00 to 18-Feb 19:30:

Data gap 17-Feb 15:00 Data gap 17-Feb 16:30 Data gap 17-Feb 19:00 Data gap 17-Feb 22:00 Data gap 18-Feb 00:00 Data gap 18-Feb 05:00 Data gap 18-Feb 06:30 Data gap 18-Feb 11:30 Data gap 18-Feb 13:00 Data gap 18-Feb 19:00


Cold Plate Temperatures 1 and 2 showed to be stable during the whole week. Cold Tip Temperatures 1 and 2 had values in accordance with the Cryocooler Power. However, because the cooler is now running at 97.6% of maximum allowed power, next Tohban should keep an eye on the Cryocooler parameters.

Detector issues

D2 is showing the same light leak signatures. In addition Rear Livetime percentages has been approximately 70%. Next Tohban should raise the threshold up. All other detectors seem stable.

Other notes

X-ray background emission is around B2. Perfect time to ask for an observation of NuSTAR.

Spacecraft Management

Decimation Normal/Vigorous
Night time data (fronts) full nighttime events
Night time data (rears) full nighttime events
Require extra passes? Yes
Requirement for moving pointer? No
Attenuator operation Normal
Personal tools